OpenStudy (anonymous):

Convert the decimal to a fraction. 0.352

OpenStudy (anonymous):

44/125

OpenStudy (anonymous):

medal plz

OpenStudy (anonymous):

how do I give a medal

OpenStudy (studygurl14):

0.352 reads 352 thousandths. That's a clue right there: \[\frac{352}{1000}\] then simplify (i'm dividing by 4 first) \[\frac{88}{250}\] then again by 2 \[\frac{44}{125}\]

OpenStudy (anonymous):

press best response

OpenStudy (studygurl14):

hey @THEUSER822 i'll give you medal if you give me one

OpenStudy (anonymous):

sure

OpenStudy (studygurl14):

awesome thx

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

thanks guys

OpenStudy (anonymous):

see and fan me

OpenStudy (anonymous):

plz fan me sorry

OpenStudy (anonymous):

i did

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

np